OpenStudy (anonymous):

Solve this DE: ydx-4(x+y^6)dy=0

OpenStudy (anonymous):

the chapter it is in is linear equations but i can't seem to figure out how

OpenStudy (anonymous):

i thought in order to do this you had to set it up as y'+p(x)y=g(x) however this is p(x,y)

OpenStudy (anonymous):

the only ways up to this point were seperation of variables and linear so....

OpenStudy (anonymous):

jim do you see a possible way?

OpenStudy (anonymous):

would this be like an implicit solution?

OpenStudy (anonymous):

\[e^{\int{}{}\frac{1}{-4(x+y^6)}dx} \]

OpenStudy (anonymous):

in which you'd make u=-4(x+y^6)=-4x-4y^6 du=-4dx du/-4=dx?

OpenStudy (anonymous):

i don't believe there was an implicit example in the book but that's the only way i can think =/

OpenStudy (anonymous):

this was in linear equations sooo i have to write it sa y' +p(x)y=g(x)=0 \[ydx-4(x+y^6)dy=0\] divide by dx \[y-4(x+y^6)y'=0\] divide both by -4(x+y^6) \[\frac{1}{-4(x+y^6)}y+y'=0=y'+\frac{1}{-4(x+y^6)}y\]

OpenStudy (anonymous):

seems i've stumped everyone .....

OpenStudy (anonymous):

I think i might have figured it out

OpenStudy (anonymous):

if i divide by dy and then distribute out

OpenStudy (anonymous):

i get \[y\frac{dx}{dy}-4x-4y^6=0\] add 4y^6 \[y\frac{dx}{dy}-4x=4y^6\] divide by y \[\frac{dx}{dy}-\frac{1}{y}x=4y^5\]

OpenStudy (anonymous):

am ion the right track by finding the reciprical

OpenStudy (anonymous):

and this answer would be implicit to the solution y'

OpenStudy (anonymous):

because this equation is linear to the variable y not x x'+p(y)x=g(y)

OpenStudy (anonymous):

i mean it's linear to the variable x not y

OpenStudy (anonymous):

\[\frac{y\cdot dx - x\cdot dy}{y}=4y^5dy \implies\frac{y\cdot d(x) - x\cdot d(y)}{y^2}=4y^4dy \]\[\implies d\left(\frac{x}{y}\right) = d\left(\frac{4y^5}{5}\right)\implies \frac{x}{y}= \frac{4y^5}{5} + C\]

OpenStudy (anonymous):

Am I right?

OpenStudy (anonymous):

not even close

OpenStudy (anonymous):

wait let me check

OpenStudy (anonymous):

because you wrote the wrong equation :/ check your last post. forgot to carry the four.

OpenStudy (anonymous):

it's gotta be linear in aspect to x so i think i have it rightly situated

OpenStudy (anonymous):

answer is \[x=2y^6+cy^4\]

OpenStudy (experimentx):

try solving for dx/dy instead ..

OpenStudy (anonymous):

solved it and got it right i got a new question with no answer tho so